# note-name = , use-sort-key = false

### v3

@@ IV.2 Steady-state time for selected initial states
Despite the nonlinear nature of the system, the change in $\Delta$ over different working regimes is intriguingly simple. Subsequent values of $\left|\mathrm{Re}\left(\lambda_{j}\right)\right|$ are packed close enough that the spectral decomposition coefficients $c_{j}$—which additionally depends on the initial state, as given in Eq. (23)—may significantly affect the system’s transient dynamics. It is thus interesting to study how fast different initial states $\rho^{\mathrm{(0)}}$ can actually reach the steady state $\rho^{\mathrm{(ss)}}$, and how different working regimes influence how fast $\rho^{\mathrm{(ss)}}$ is reached for the given $\rho^{\mathrm{(0)}}$.
@@
One measure for the transient lifetime is the relaxation time $\tau$, which measures the time at which the difference between some observable and its steady-state value reaches $1/e$ the value at the beginning of the evolution [83]. We are, however, interested in the actual time it takes for the system to reach $\rho^{\mathrm{(ss)}}$. Adding to this the nonlinear nature of the system, we instead use the steady-state time $T_{\mathrm{ss}}$, the criterion for which is given by Eq. (14).
@@
Obviously, we cannot study the evolution of all density matrices. Here we consider three kinds of initial states. The first is the Fock state $\rho_{\mathrm{Fock}}=\left|{n}\right\rangle\left\langle{n}\right|$, which we choose for its simplicity as an eigenstate of the harmonic oscillator that exhibits Wigner negativity. Next, we consider the thermal state $\rho_{\mathrm{th}}=\sum_{k}\left[\tilde{n}_{\mathrm{th}}^{k}/\left(1+\tilde{n}_{\mathrm{th}}\right)^{k+1}\right]\left|{k}\right\rangle\left\langle{k}\right|$ where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=\tilde{n}_{\mathrm{th}}$. With this state, we study how the system’s transient dynamics behave starting in a thermal equilibrium. Lastly, we study the coherent state $\rho_{\mathrm{coh}}=\left|{\beta}\right\rangle\left\langle{\beta}\right|=\exp\left(-|\beta|^{2}\right)\sum_{mn}\left(\beta^{m}\beta^{*n}/\sqrt{m!n!}\right)\left|{m}\right\rangle\left\langle{n}\right|$ where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=|\beta|^{2}$, which possesses quantum coherence. To compare different initial states, we start with the same energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ so the early dynamics is only influenced by the $\gamma_{2}$ process, as evident from Eq. (8). This is as far as we can get to putting them on equal footing. We choose $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}=1,3,10$ for each initial state, which we deem sufficient to show the change in the behavior of $T_{\mathrm{ss}}$.
@@
For the Fock and thermal states, the behavior of $T_{\mathrm{ss}}$ shows a feature that is absent in $\Delta$: over some parts of the system’s working regimes, $T_{\mathrm{ss}}$ drops to a minimum whose value is significantly smaller than the values at nearby working regime points. Additionally, when we plot $T_{\mathrm{ss}}$ against $\mathcal{B}$ for a fixed $\mathcal{A}$, we see a maximum for $T_{\mathrm{ss}}$ which occurs at a value of $\mathcal{B}$ larger than that at which the minimum occurs. This is easier to see in thermal states for our choice of initial states, but Fock states also show this behavior. Whether or not a minimum occurs for a larger $\mathcal{B}$ is beyond our knowledge. From our results, we see that this region of speedy parameters moves toward larger $\mathcal{A}$ as the initial energy is increased. Furthermore, for Fock state with $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}=10$, another region emerges in the small $\mathcal{A}$ regime. Our findings may not capture all the properties of the steady-state regions and thus provide no insight into a pattern for their occurrence, even for a given kind of initial state.
@@
Nevertheless, we can logically relate the existence of the speedy regions to the system’s energy dynamics, since the states exhibiting the speedy regions are diagonal. We note that the steady-state energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ is independent of the individual values of the nonconservative parameters, as we mentioned in Section II.5. A stronger energy gain than the nonlinear dissipation will shift $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ toward a higher energy level, where the nonlinear dissipation is stronger. As such, despite the strong energy gain, the system struggles to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. Conversely, strong nonlinearity brings $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ down where its effect becomes weaker. Despite the strong nonlinearity, the system also struggles to reach the low $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. The speedy parameters thus give the right balance for the equilibrium to be more easily attained. Additionally, where the speedy parameters occur must also depend on the initial energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$. For example, we may consider cases where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}>\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ where the energy gain has to push the system’s energy up against the increasingly strong nonlinear dissipation. It will take longer to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ the lower $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ is. This is likewise true when $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}<\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$, in which case the progressively weakened nonlinear dissipation has to cover more energy shift opposing the energy gain to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$.
@@
We find that the value of $T_{\mathrm{ss}}$ for a coherent state $\rho_{\mathrm{coh}}=\left|{\beta}\right\rangle\left\langle{\beta}\right|$ only depends on $|\beta|$ (see also the Supplementary Materials). This is not surprising. The magnitude of the coherence is the same for coherent states with the same energy: $\left|\left\langle{m}\middle|{\rho_{\mathrm{coh}}}\middle|{n\neq m}\right\rangle\right|=\exp\left(-\left|\beta\right|^{2}\right)\left|\beta\right|^{m}\left|\beta\right|^{n}/\sqrt{m!n!}$. Additionally, in the phase space representation, we can imagine ourselves on the Wigner function peak and undergoing the dynamics in the simple-harmonic rotating frame—from this perspective, the Wigner function evolves the same way regardless of the initial angular position of the peak.
Unlike the other two states, $T_{\mathrm{ss}}$ for a coherent state behaves similarly to the Liouvilian gap, as clearly shown in Fig. 8, which suggests that the $j=1$ term is indeed the last to vanish for a coherent state. As shown by the middle plot in Fig. 8, $T_{\mathrm{ss}}$ also has a local minimum corresponding to the local maximum of the Liouvillian gap. However, this is not an appreciable speedup like in the cases of the other two initial states. In Section III.2, we observe that the steady-state time for the coherence state is determined by the decay of its neighboring-level coherence. This strengthens our description of the speedy parameters in terms of the energy gain: we do not see them for a coherent state since $T_{\mathrm{ss}}$ is determined by the coherence decay instead of the energy dynamics.
@@ V Conclusions and Remarks
Like its classical counterpart, the quantum Stuart-Landau oscillator is a paradigm in the study of the quantum limit cycle and synchronization. Studies concerning its steady state have already been adequately done, revealing many interesting features despite its simplicity. To complement that, our study of the transient dynamics—how the system is attracted toward the limit cycle—has also revealed interesting transient behaviors of the system.
We have described the energy dynamics of the system and divided the transient dynamics into working regimes based on the energy balance. In particular, we have noted the linear energy gain that is not present in the classical counterpart. Whether or not this contribution is significant determines whether the system can evolve in the classical regime, where the quantum fuzziness in the phase-space representation can be interpreted. This allows us to interpret the quantum phase-space representation as a true probability distribution, exhibiting how, at the steady state, the phase point of the system can be anywhere on the classical limit cycle. Through our numerical results, we have also inferred that the system’s steady-state energy depends on the relative values between the nonconservative parameters, not their individual values.
We have studied the system’s transient phase-space dynamics. We have obtained the Wigner function equation of motion, which is similar to the Kramers-Moyal equation for stochastic processes but does not obey some of its properties due to the Wigner function being a quasiprobability distribution. Unlike the other processes, the two-quantum dissipation corresponds to a jumpy Wigner function dynamics. We have also described how the nonlinear process is necessary for the emergence of a limit cycle in terms of how different processes affect the Wigner function evolution. We have studied the classical-like behavior embodied by the coherent state. We have described how the Wigner function evolution differs in the working regime that allows a classical limit cycle, and how the coherent state takes a long time to reach the limit cycle due to the presence of quantum coherence between neighboring energy levels. As for states with quantum features, we have studied how, for some initial states, the nonlinear process can temporarily hold the quantumness of the system early in the evolution, as indicated by the temporary increase in the Wigner function’s negative volume.
Lastly, we have studied the evolution speed of the system. We have numerically calculated the eigenvalues of the system’s Lindbladian. In particular, we have found that the Liouviilian gap, which characterizes the slowest decay toward the steady state, decreases with a stronger energy gain. We have also described the general behavior of the eigenvalues and analyzed the numerical artifact that comes with the use of a truncated Hilbert space dimension. We have characterized the time it takes the Fock, thermal, and coherent states to get sufficiently close to the steady state. We have found that the dependence of the steady-state time on the system’s parameters may differ from that of the Liouvillian gap, depending on the initial states. In particular, for diagonal states, we have observed regions over the system’s working regimes for which the steady-state time is significantly reduced. We have attributed the absence of the speedy parameters for coherent states to the slow coherence decay.
@@
Our findings contribute to a further understanding of the quantum limit cycle attraction. Nevertheless, some questions remain. We have seen that the coherence of a coherent state holds it back from reaching the steady state, leading to a steady-state time that is magnitudes larger than other states such as the Fock state and the thermal state. More generally, quantum coherence slowing down the limit cycle attraction may not be desirable in some scenarios, such as when we are considering the limit cycle’s robustness against coherence-generating interactions. Is there any useful way to speed up coherence decay while keeping the limit cycle unchanged? Additionally, we have seen that the speedy parameters offer a significant reduction in steady-state time. Is there a general pattern for their occurrence beyond what we have discussed? Moreover, do they result in a speedy synchronization? Even if they do not, it is interesting to see whether similar speedy parameters exist when the system is coupled to one another. Additionally, it is interesting to see whether the speedy parameters can occur for nondiagonal states. This may be possible if the quantum coherence can decay faster than the time it takes for the energy to reach its steady-state value, e.g. for a cat state with many cats.
@@ Acknowledgements.
We thank Mahameru BRIN for their HPC facilities. We acknowledge the use of the following numerical packages: NumPy, SciPy, SymPy, QuTiP, and Matplotlib. H.M.L. is supported by research assistantship from BRIN Directorate for Talent Management. The data supporting the findings of this study are available from the corresponding authors upon reasonable request. The Python package for Appendix B is publicly available at https://github.com/hendry24/moyalstar.git.
@@ Appendix A Wigner function and the system’s energy
$\mathrm{tr}\left({\hat{\mathcal{F}}\hat{\mathcal{G}}}\right)=2\pi\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\hat{\mathcal{F}}}(x,p)W_{\hat{\mathcal{G}}}(x,p)\ \mathrm{d}x\ \mathrm{d}p$ (26)
@@
$\hat{\mathcal{F}}$ $=\rho=\sum_{\ell}\rho_{\ell\ell}\left|{\ell}\right\rangle\left\langle{\ell}\right|+\sum_{m,n\neq m}\rho_{mn}\left|{m}\right\rangle\left\langle{n}\right|$ (27a)
$\hat{\mathcal{G}}$ $=\hat{a}^{\dagger}\hat{a}=\sum_{k}k\left|{k}\right\rangle\left\langle{k}\right|$ (27b)
@@
$\begin{split}W(x,p)&=W_{\mathrm{D}}(x,p)+W_{\mathrm{OD}}(x,p)\\ &=\sum_{\ell}\rho_{\ell\ell}W_{\left|{\ell}\right\rangle\left\langle{\ell}\right|}(x,p)+\sum_{m,n\neq m}\rho_{mn}W_{\left|{m}\right\rangle\left\langle{n}\right|}(x,p)\end{split}$ (28)
@@
$W_{\hat{a}^{\dagger}\hat{a}}(x,p)=\sum_{k}kW_{\left|{k}\right\rangle\left\langle{k}\right|}(x,p)$ (29)
@@
$\begin{split}2\pi&\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\mathrm{OD}}(x,p)W_{\hat{a}^{\dagger}\hat{a}}(x,p)\ \mathrm{d}x\ \mathrm{d}p\\ &=\sum_{k,m,n\neq{m}}k\rho_{mn}\mathrm{tr}\left({\left|{k}\right\rangle\left\langle{k}\middle|{m}\right\rangle\left\langle{n}\right|}\right)\\ &=0\end{split}$ (30)
@@
$\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=2\pi\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\mathrm{D}}(x,p)W_{\hat{a}^{\dagger}\hat{a}}(x,p)\ \mathrm{d}x\ \mathrm{d}p$ (31)
@@
It is straightforward to generalize this result to any expectation value that depends only on the diagonal elements, such as the energy variance—such a quantity is embodied by only the rotationally invariant part of the system’s Wigner function.

### v5

@@ IV.2 Steady-state times for selected initial states
Despite the nonlinear nature of the system, the change in $\Delta$ over different working regimes is intriguingly simple. Subsequent values of $\left|\mathrm{Re}\left(\lambda_{j}\right)\right|$ are packed close enough that the spectral decomposition coefficients $c_{j}$, which additionally depends on the initial state, as given in Eq. (23), may significantly affect the system’s transient dynamics. It is thus interesting to study how fast different initial states $\rho^{\mathrm{(0)}}$ can actually reach the steady state $\rho^{\mathrm{(ss)}}$, and how different working regimes influence how fast $\rho^{\mathrm{(ss)}}$ is reached for the given $\rho^{\mathrm{(0)}}$.
@@
One measure for the transient lifetime is the relaxation time $\tau$, which measures the time at which the difference between some observable and its steady-state value reaches $1/e$ of the value at the beginning of the evolution [85]. We are, however, interested in the actual time it takes for the system to reach $\rho^{\mathrm{(ss)}}$. Adding to this the nonlinear nature of the system, we instead use the steady-state time $T_{\mathrm{ss}}$, the criterion for which is given by Eq. (14).
Obviously, we cannot study the evolution of all density matrices. Here we consider three kinds of initial states. The first is the Fock state
$\rho_{\mathrm{Fock}}=\left|{n}\right\rangle\left\langle{n}\right|,$ (26)
which we choose for its simplicity as an eigenstate of the harmonic oscillator that exhibits Wigner negativity. Next, we consider the thermal state
$\rho_{\mathrm{th}}=\sum_{k}\frac{\tilde{n}_{\mathrm{th}}^{k}}{\left(1+\tilde{n}_{\mathrm{th}}\right)^{k+1}}\left|{k}\right\rangle\left\langle{k}\right|,$ (27)
to study the system’s transient limit cycle attraction starting in a thermal equilibrium with an average energy of $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=\tilde{n}_{\mathrm{th}}$. Lastly, we study the coherent state, whose density matrix is given by
$\rho_{\mathrm{coh}}=\left|{\beta}\right\rangle\left\langle{\beta}\right|=e^{-|\beta|^{2}}\sum_{mn}\frac{\beta^{m}\beta^{*n}}{\sqrt{m!n!}}\left|{m}\right\rangle\left\langle{n}\right|,$ (28)
@@
where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=|\beta|^{2}$, which possesses quantum coherence. To compare different initial states, we start with the same energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ so the early dynamics is only influenced by the $\gamma_{2}$ process, as evident from Eq. (8). This is as far as we can get to putting them on equal footing. We choose $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}=1,3,10$ for each initial state, which we deem sufficient to show the change in the behavior of $T_{\mathrm{ss}}$.
@@
For the Fock and thermal states, the behavior of $T_{\mathrm{ss}}$ shows a feature that is absent in $\Delta$: over some parts of the system’s working regimes, $T_{\mathrm{ss}}$ drops to a minimum whose value is significantly smaller than the values at nearby working regime points. Additionally, when we plot $T_{\mathrm{ss}}$ against $\mathcal{B}$ for a fixed $\mathcal{A}$, we see a maximum for $T_{\mathrm{ss}}$ which occurs at a value of $\mathcal{B}$ larger than that at which the minimum occurs. This is easier to see in thermal states for our choice of initial states, but Fock states also show this behavior. Whether or not a minimum occurs for a larger $\mathcal{B}$ is beyond our knowledge. From our results, we see that this region of fast convergence regimes moves toward larger $\mathcal{A}$ as the initial energy is increased. Furthermore, for Fock state with $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}=10$, another region emerges in the small $\mathcal{A}$ regime. Our findings may not capture all the properties of the steady-state regions and thus provide no insight into a pattern for their occurrence, even for a given kind of initial state.
@@
Nevertheless, we can logically relate the existence of the speedy regions to the system’s energy dynamics, since the states exhibiting the speedy regions are diagonal. We note that the steady-state energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ is independent of the individual values of the nonconservative parameters, as we mentioned in Section II.5. A stronger energy gain than the nonlinear dissipation will shift $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ toward a higher energy level, where the nonlinear dissipation is stronger. As such, despite the strong energy gain, the system struggles to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. Conversely, strong nonlinearity brings $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ down where its effect becomes weaker. Despite the strong nonlinearity, the system also struggles to reach the low $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. The fast convergence regimes thus give the right balance for the equilibrium to be more easily attained. Additionally, where the fast convergence regimes occur must also depend on the initial energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$. For example, we may consider cases where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}>\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ where the energy gain has to push the system’s energy up against the increasingly strong nonlinear dissipation. It will take longer to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ the lower $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$ is. This is likewise true when $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}<\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{(0)}$, in which case the progressively weakened nonlinear dissipation has to cover more energy shift opposing the energy gain to reach $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$.
We find that the value of $T_{\mathrm{ss}}$ for a coherent state $\rho_{\mathrm{coh}}=\left|{\beta}\right\rangle\left\langle{\beta}\right|$ only depends on $|\beta|$. This is not surprising: the magnitude of the coherence is the same for coherent states with the same energy, i.e.,
$\left|\left\langle{m}\middle|{\rho_{\mathrm{coh}}}\middle|{n\neq m}\right\rangle\right|=e^{-\left|\beta\right|^{2}}\frac{\left|\beta\right|^{m}\left|\beta\right|^{n}}{\sqrt{m!n!}}.$ (29)
Additionally, in the phase space representation, we can imagine ourselves on the Wigner function peak and undergoing the dynamics in the simple-harmonic rotating frame. From this perspective, the Wigner function evolves the same way regardless of the initial angular position of the peak.
Unlike the other two states, $T_{\mathrm{ss}}$ for a coherent state behaves similarly to the Liouvilian gap, as clearly shown in Fig. 8, which suggests that the $j=1$ term is indeed the last to vanish for a coherent state. As shown by the middle plot in Fig. 8, $T_{\mathrm{ss}}$ also has a local minimum corresponding to the local maximum of the Liouvillian gap. However, this is not an appreciable speedup like in the cases of the other two initial states. In Section III.2, we observe that the steady-state time for the coherence state is determined by the decay of its neighboring-level coherence. This strengthens our description of the fast convergence regimes in terms of the energy gain: we do not see them for a coherent state since $T_{\mathrm{ss}}$ is determined by the coherence decay instead of the energy dynamics.
@@ V Conclusions and Remarks
The quantum Stuart-Landau oscillator is a paradigm in the study of quantum limit cycle and synchronization. We have studied its transient dynamics and observed intriguing behaviors. We have formulated the criteria for the system to enter its classical regime, where the steady-state Wigner function is localized near its circular peak, minimizing the quantum noise. We have evaluated the equation of motion for the Wigner function and compared it to the Kramers-Moyal equation for stochastic processes. It demonstrates how the quasiprobabilistic nature of the Wigner function enables the contribution from the jumpy two-quantum dissipation to disregard the Pawula theorem, and also how this process is essential for the limit cycle to form. We have shown that this process is also responsible for the temporary increase of Wigner negativity. We have compared the evolution of a coherent state inside and outside the system’s classical regime, and observed how the evolution is slow due to the slow decay of neighboring-level coherence. We have numerically evaluated the system’s Lindbladian eigenvalues and observed the decrease in Liouvillian gap with stronger energy gain. Evolving the system from selected initial states, we have found that the dependence of the time to converge toward the limit cycle does not generally follow the trend of the Liouvillian gap. In particular, we have observed fast convergence regimes in the system’s parameter space, over which the steady-state time is significantly reduced locally; these are not observed for the coherent states due to the lingering coherence.
@@
These findings provide deeper insights into the transient behavior of the quantum Stuart-Landau oscillator and a better understanding of self-sustained quantum systems in general. Nevertheless, some questions remain. We have seen that the coherence of a coherent state holds it back from reaching the steady state, leading to a steady-state time that is magnitudes larger than that of other states, such as the Fock state and the thermal state. More generally, quantum coherence slowing down the limit cycle attraction may not be desirable in some scenarios, such as when we are considering the limit cycle’s robustness against coherence-generating interactions. Is there any useful way to speed up coherence decay while keeping the limit cycle unchanged? Additionally, we have seen that the fast convergence regimes offer a significant reduction in steady-state time. Is there a general pattern for their occurrence beyond what we have discussed? Moreover, do they result in a speedy synchronization? Even if they do not, it is interesting to see whether similar regimes exist when the system is coupled with another quantum self-sustained oscillator. Additionally, we are curious whether the fast convergence regimes can occur for nondiagonal states. This may be possible if the quantum coherence can decay faster than the time it takes for the energy to reach its steady-state value, e.g., for a cat state with many cats.
@@ Acknowledgements.
We thank Mahameru BRIN for their HPC facilities. We acknowledge the use of the following numerical packages: NumPy [87], SciPy [88], SymPy [89], QuTiP [90, 91], Matplotlib [92], and moyalstar [93]. H.M.L. is supported by a research assistantship from the BRIN Directorate for Talent Management. The data supporting the findings of this study are available from the corresponding authors upon reasonable request.
@@ Appendix A Wigner function and the system’s energy
$\mathrm{tr}\left({\hat{\mathcal{F}}\hat{\mathcal{G}}}\right)=2\pi\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\hat{\mathcal{F}}}(x,p)W_{\hat{\mathcal{G}}}(x,p)\ \mathrm{d}x\ \mathrm{d}p,$ (30)
@@
$\hat{\mathcal{F}}$ $=\rho=\sum_{\ell}\rho_{\ell\ell}\left|{\ell}\right\rangle\left\langle{\ell}\right|+\sum_{m,n\neq m}\rho_{mn}\left|{m}\right\rangle\left\langle{n}\right|,$ (31a)
$\hat{\mathcal{G}}$ $=\hat{a}^{\dagger}\hat{a}=\sum_{k}k\left|{k}\right\rangle\left\langle{k}\right|,$ (31b)
@@
$\begin{split}W(x,p)&=W_{\mathrm{D}}(x,p)+W_{\mathrm{OD}}(x,p)\\ &=\sum_{\ell}\rho_{\ell\ell}W_{\left|{\ell}\right\rangle\left\langle{\ell}\right|}(x,p)+\sum_{m,n\neq m}\rho_{mn}W_{\left|{m}\right\rangle\left\langle{n}\right|}(x,p).\end{split}$ (32)
@@
$W_{\hat{a}^{\dagger}\hat{a}}(x,p)=\sum_{k}kW_{\left|{k}\right\rangle\left\langle{k}\right|}(x,p),$ (33)
@@
$\begin{split}0&=2\pi\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\mathrm{OD}}(x,p)W_{\hat{a}^{\dagger}\hat{a}}(x,p)\ \mathrm{d}x\ \mathrm{d}p\\ &=\sum_{k,m,n\neq{m}}k\rho_{mn}\mathrm{tr}\left({\left|{k}\right\rangle\left\langle{k}\middle|{m}\right\rangle\left\langle{n}\right|}\right).\end{split}$ (34)
@@
$\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=2\pi\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}W_{\mathrm{D}}(x,p)W_{\hat{a}^{\dagger}\hat{a}}(x,p)\ \mathrm{d}x\ \mathrm{d}p.$ (35)
@@
It is straightforward to generalize this result to any expectation value that depends only on the diagonal elements, e.g., the energy variance. Such a quantity is embodied by only the rotationally invariant part of the system’s Wigner function.